Five computers will be delivered to a certain school for testing, but three of them are defective. The principal will get two of these computers for his office. Let T represent the random variable representing the defective computers will occur. Find the values of random variable T.  



1
Expert's answer
2021-02-04T08:12:31-0500

Given that out of "5" computers 3 computers are defective.

The principal will get "2" of those.

Out of "5" computers "2" can be choose in "^5C_2" ways.

Therefore sample space "S" contains "^5C_2" number of points.

Out of "5" computers principal will get "2" computers in such way

(i) "2" computers are defective and it can be choose in "^3C_2" ways.

(ii) "1" computer is defective and "1" computer is good, it can be choose in "^2C_1\u00d7^3C_1" ways.

(iii) "2" computers are good and it can be choose in "^2C_2" ways.

Let "T" be a random variable representing number of defective computers.

Then "T" can take the values "0,1,2."

Now we will find the probabilities of each points corresponding to the random variable "T."

"\therefore P(T=0)=" probability of getting no defective computers

"=\frac{^2C_2}{^5C_2}=\frac{1}{10}"

"P(T=1)=" probability of getting "1" defective computer

"=\frac{^2C_1×^3C_1}{^5C_2}=\frac{6}{10}=\frac{3}{5}"

"P(T=2)=" probability of getting "2" defective computers

"=\frac{^3C_2}{^5C_2}=\frac{3}{10}"

Which are the required values of random variable "T."